A Covid 19 testing machine at a testing centre breaks down an average of three times per year. Using an appropriate probability distribution formula, find the probability that during the next year, this machine will have;

i. exactly two breakdowns

ii. at most one breakdown


1
Expert's answer
2020-07-07T20:06:43-0400

solution:


The number of successes in two disjoint time intervals is independent. The probability of a success during a small time interval is proportional to the entire length of the time interval


given : "\lambda" = 3


Formula: P(X=x)= "\frac{\lambda ^{x}e^{\lambda }}{x!}"


part i)

Exactly two breakdowns

P(x=2)= "\frac{3 ^{2}e^{3 }}{2!}" = 0.2240


Part ii)

At most one breakdown:

P(X"\leq" 1)= P(X=0)+P(x=1)

="\frac{3 ^{0}e^{3 }}{0!}" + "\frac{3 ^{1}e^{3 }}{1!}"

=0.0498+0.1494

=0.1992